Constants 1 Periodic Table Part A Two small nonconducting spheres have a total charge of 95.0 μC When placed 1.11 m apart, the force each exerts on the other is 11.0 N and is repulsve. What is the charge on each? Enter your answers in ascending order separated by comma. 1-2.01 10$7.49M10s Correct ▼PartB What if the force were attractive? Enter your answers in ascending order separated by comma. 44.-1.09 .10 4. 1.4-10 Submit Previows Answers Request Ansr X Incorrect: Try Again

Please answer part 2 thoroughly, I do not understand how my answer is wrong.
Answer #1

Part B)

Qi + Q2 = 95 μC = 95 * 10-6 C

Force between the two charges is attractive, so if 0 is positive Q_2 must be negative.

F=-11.0N ( negative sign indicates the force is attractive)

TEQr

8.9910 10211.0

Q2-一1.05076 * 10-9

Q_1-Q_2=sqrt{(Q_1+Q_2)^2-4Q_1Q_2}

Q_1-Q_2=sqrt{(95*10^{-6})^2-4(-1.5076*10^{-9})}=0.0001227

Qi-Q2 0.0001227 ---(1)

91 + Q2 0.000095 ---(2)

adding equations (1) and (2) 201 = 0.0002177 Qi = 0.000109 ,

using the above value in (1) 0.000 109-Q2 = 0.000 1227 Q2ー-0.000014

  Q_1=1.09*10^{-4},C and 21.410-5 C

(Verification:

Qi + Q2 = 0.000 109 + (-0.000014) = 0.000095 = 95 * 10-6 C)